3. On 1/1/2015, Choco paid $92,000 to acquire 10% of the voting common stock of Cookie. In 1/1/2016, Choco acquired additional 20% of the voting common stock of Cookie for $210,000. Following is the financial information about Cookie. Book value of net assets 1/1/2015 $800,000 Net income (2015) Net income (2016) $180,000 $210,000 Dividends (2015) Dividends (2016) $80,000 $100,000 Land undervalued 12/31/2015 Land undervalued 12/31/2016 $60,000 $70,000 All excess payment will be recorded using Trademark which has useful life of 9 years in 2016. During the year, there was no fair market value adjustment for Cookie.

Q9: Calculate annual amortization of under/overvalued asset and Trademark in 2016.

Q10: In 2016, what is the balance of the investment account in Cookie at 12/31/2016 using the equity method?
Answer #1

Q 9: Annual amortization of under/overvalued asset and Trademark in 2016:

Since Land has been under valued and it is a non- depreciable asset, hence no amortization will be recorded for the same in books of accounts.

With regard to Trademark, the annual amortization will be of amount spread over life of 9 years. However, every year it will be tested for certainity of recognition.

Q 10: Balance of Investment account in 2016

Particulars 2015 2016
Investment made by Choco in Cookie 92,000 (10%) 210,000 (additional 20%)
Net income of Cookie 180,000 210,000
Dividend paid by Cookie 80,000 100,000

Value of investment made by Choco will be reduced by dividend and net income which it has earned from Cookie as it is an income inflow to Choco for investment made.

In 2015 = 92,000 - (180,000+80,000)*10% = $66,000

In 2016= Investment received - Dividend paid - Net income attributable to Choco Investment

i.e. (66,000 + 210,000) - (100,000+210,000)*30% = $183,000